I have been having problems solving for x in the following equation.
3x^2+10x+3/x+3=x^2-4x+3/x-1
Thanks Ian

Answer by apachemexican1(64)

okay,
what you wanna do first is,
fix the equation to where 3x^ and x^2 are on one side and the rest
of the equation is on the other
3x^2-x^2=10x-3/x-3-4x+3/x-1
then combine like terms
2x^2=-14x-4 (-3/x+3/x=0; they cancel each other out.)
then subtract
-2x from both sides
x=-16-4
then subtract again
x=-20